Assume the nominal rate of return is 8.63% and the real rate is 4.68%. Find the inflation rate of return using the exact formula.

Answer Format: Percentage Round to: 0 decimal places (Example: 9%, % sign required. Will accept decimal format rounded to 2 decimal places (ex: 0.09))
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.0468)=(1+0.0863)/(1+inflation rate)

1+inflation rate=(1.0863/1.0468)

inflation rate=(1.0863/1.0468)-1

which is equal to

=4%(Approx).